OpenStudy (anonymous):

@iGreen. A drinking glass is shaped like a cylinder with a height of 5 in. and a diameter of 3 in. Maeva adds 25 spherical pieces of ice to the glass. The pieces of ice each have a diameter of 1.1 in. How many cubic inches of water can Maeva add to fill the glass to the rim? Use 3.14 to approximate pi and round your answer to one decimal place. ___ in^3

OpenStudy (anonymous):

Okay, we have to subtract the volume of the ice pieces from the volume of the cylinder glass.

OpenStudy (anonymous):

Volume of a cylinder: \(V = \pi r^2 h\) Plug in what we know: \(V = (3.14)(1.5^2)(5)\) Simplify

OpenStudy (anonymous):

1.5 times 1.5 = 2.25

OpenStudy (anonymous):

3.14 times 2.25 times 5

OpenStudy (anonymous):

35.325

OpenStudy (anonymous):

is that my answer? 35.325? or would i simplify more?

OpenStudy (anonymous):

No, now we have to find the volume of each spherical piece of ice: \(V = \dfrac{4}{3} \pi r^3\) Plug in what we know: \(V = \dfrac{4}{3} (3.14) (0.55)^3\) Simplify

OpenStudy (anonymous):

0.55^3 = 0.166375

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

1.3333 times 3.14 times 0.166375

OpenStudy (anonymous):

Yes^^

OpenStudy (anonymous):

Multiply them

OpenStudy (anonymous):

0.69653925275

OpenStudy (anonymous):

Yes, that is the volume for each piece, and there are 25 pieces, multiply by 25.

OpenStudy (anonymous):

17.4134813187

OpenStudy (anonymous):

Yes, now subtract: 35.325 - 17.4134813187

OpenStudy (anonymous):

That will be your answer.

OpenStudy (anonymous):

17.9115186813

OpenStudy (anonymous):

(It must be a decimal and consist of 3 numbers) So 17.9?

OpenStudy (anonymous):

Yes, correct.

OpenStudy (anonymous):

Alright cool. Thanks! :)

OpenStudy (anonymous):

Np